OpenStudy (dr0zier99):

WILL FAN AND MEDAL Which equation is in point-slope form and depicts the equation of this line?

OpenStudy (mayankdevnani):

use two point slope form

OpenStudy (anonymous):

y-y*1 = m(x-x*1)

OpenStudy (mayankdevnani):

http://www.purplemath.com/modules/strtlneq2.htm

OpenStudy (mayankdevnani):

Hope it helps ! @dr0zier99

OpenStudy (dr0zier99):

the answerws are (y+4)=-1/3(x+1) (y-1)=-1/3(x-2) (y-4)=-5/3(x-1) (y+4)=5/3(x+1)

OpenStudy (mayankdevnani):

what do you think ?

OpenStudy (dr0zier99):

idk im kinda comfused

OpenStudy (dr0zier99):

from what the purple math says i would think that it is c?

OpenStudy (mayankdevnani):

 \[\large \bf slope~form=>y-y_1=m(x-x_1)\] \[\large \bf ~where~m=slope=\frac{y_2-y_1}{x_2-x_1}\]